OpenStudy (angel_kitty12):

The students of a high school are 51% males; 45% of the males and 49% of the females attend concerts. A. Find the probability that a student attends concerts. B. Find the probability that a student is a female and does not attend concerts.

OpenStudy (anonymous):

a. is 1/2

OpenStudy (anonymous):

b. 1/2 x 49/100 x 49/100

OpenStudy (angel_kitty12):

How is a 1/2 though?

jimthompson5910 (jim_thompson5910):

Let's say there are 10,000 students 51% are males, so 0.51*10,000 = 5,100 are male 49% are female, so 0.49*10,000 = 4,900 are female ------------------------------------------------------- `45% of the males ... attend concerts` 0.45*5100 = 2,295 males attend concerts `49% of the females attend concerts` 0.49*4900 = 2,401 females attend concerts In total 2295+2401 = 4,696 students attend concerts

jimthompson5910 (jim_thompson5910):

We have 4,696 students who attend concerts out of 10,000 total

jimthompson5910 (jim_thompson5910):

probability student attends concert = (# of students who attend concert)/(# total)

OpenStudy (angel_kitty12):

Thank you :)

jimthompson5910 (jim_thompson5910):

no problem